OpenStudy (queen2988):

76 is .2% of ______.

OpenStudy (anonymous):

well how do we find the percent of a number

OpenStudy (xapproachesinfinity):

it goes like this 76=0.2% x=100% cross multiply 0.2x=76.100 x=7600/.2

OpenStudy (xapproachesinfinity):

is that 2% or .2% lol

OpenStudy (xapproachesinfinity):

not sure what he wrote .2 or 2? hehe

OpenStudy (xapproachesinfinity):

either way, the what is important is that he learns how to get it hehe

OpenStudy (anonymous):

I would use a formula, is/of equals percent/100. Just plug everything in you should find your answer.

OpenStudy (skullpatrol):

$$\Huge \frac{x}{100}=\frac{a}{b}$$ $$\Huge \frac{0.2}{100}=\frac{76}{b}$$

OpenStudy (xapproachesinfinity):

@smallslies use what he said more easier 76=0.2%x x=(76.100)/.2 .2 percent is the same as (.2/100) is means equal, of mean times (product)

OpenStudy (anonymous):

@xapproachesinfinity Alright, it would be very confusing to understand my explanation. thanks for seeing that

OpenStudy (xapproachesinfinity):

@smallslies, there are many ways to this but the result is one hehe